# Implicit Neural Representations for Spatial Transcriptomics

### v1

@@ 3.2.4 Decoding Head
A straightforward approach to decoding $\hat{\mathbf{z}}$ back to raw representations is to directly use the pre-trained GAE decoder for end-to-end INR tuning. However, this method encounters two primary challenges. (1) The embedding mapping of $\mathbf{x}\to\mathbf{z}$ is error-prone, and domain shifts can severely impair decoding performance, compounding with the inherent imperfections in GAE’s reconstruction. (2) A pre-trained decoder can become trapped in local minima, impeding the optimization of the embedding mapping if the INR depends on gradients from the decoder.
To address (1)(2), we first warm up the INR with Eq. 2 alone to stabilize the mapping, then attach the pre-trained decoder to learn the mapping from $\mathbf{z}\to\mathbf{y}$, keeping the INR fixed. This decoder-only training phase is designed to finetune a case-specific decoder that compensates for mapping errors and minimizes cumulative errors.
In INR regression tasks, norm-based loss functions (e.g., $\ell^{2}$ norm or MSE) are typically used, as they assume normally (Gaussian) distributed errors. However, for zero-inflated ST data, this assumption is invalid. To handle the imbalanced distribution of zero and non-zero values, we apply Dice Loss [45], which is sensitive to class imbalance and treats the regression task as a quasi-classification one. Dice Loss optimizes for Intersection over Union (IoU) between the prediction map and binary ground truth. Specifically, we use the non-negative half of $\tanh(\cdot)$ to map network outputs into a pseudo probability into a pseudo-probability range $[0,1)$, and compute the intersection using element-wise Hadamard products with the ground truth. To avoid division by 0, Dice Loss is computed as:
@@ 4.1 Datasets & Metrics
For a quantitative benchmarking, we involve a nanoscale resolution Stereo-seq (SpaTial Enhanced REsolution Omics-sequencing) dataset, MOSTA [2]. MOSTA consists of a total of 53 sagittal sections from C57BL/6 mouse embryos at 8 progressive stages using Stereo-seq, from which we take 1 slice for each stage (from E9.5 to E16.5) for benchmarking. In addition to Stereo-seq, we also leverage ST data by two other common platforms, including Slide-seqV2 and 10x Genomics Visium, to demonstrate the generalization of SUICA.
As for evaluation of the fitting performance of SUICA with super-high dimensional data, we focus on 3 aspects, namely numerical fidelity, statistical correlation, and bio-conservation. For numerical fidelity, we apply MSE, MAE (Mean Absolute Error) and cosine similarity to measure the significant, subtle and directional errors between the predicted and ground-truth values respectively. Note that we only measure numerical fidelity on non-zero values considering the zero-inflated distribution of ST. To measure the statistical correlation, we employ Pearson Correlation Coefficient and Spearman’s Rank Correlation Coefficient (Spearman’s $\rho$), with both of them ranging from -1 to 1. Lastly, to evaluate how well the prediction preserves cellular heterogeneity and spatial coherence within the microenvironment of the slice, we use the Adjusted Rand Index (ARI) as a metric for quantifying bio-conservation leveraging the independent hand-crafted cell type annotations.
@@ 4.2 Implemented Details
We implement SUICA with PyTorch, sampling 80% of spots within a ST slice for training and retaining the remaining 20% for evaluation, under the evaluation protocol of super-resolution. Notably, this sampling scheme may place some validation spots outside the convex hull of the training set, making such interpolation challenging. To construct the $k$-NN graph for GCN, we set $k=5$, including the given cell itself. During the pre-training phase of GAE, we use Adam with a learning rate of $1e^{-5}$ for 200 epochs. After obtaining low-dimensional cell embeddings, we train the INR with Adam at a learning rate of $1e^{-4}$ for $1k$ epochs to learn the embedding mapping. Subsequently, the INR is frozen, with the pretrained decoder trained for an additional $1k$ epochs using Adam with the same learning rate. The deep learning-based reference-free cell super-resolution model, STAGE [21], is trained till convergence for benchmarking. All experiments are conducted on 1 NVIDIA RTX 4090.
4.3 Benchmarking Results
@@
Methods Numerical Fidelity Statistical Correlation Bio-conservation
@@
Nearest Interp. 6.24 1.11 0.741 0.695 0.353 0.280
FFN [48] 6.51 1.20 0.706 0.718 0.400 0.143
SIREN [43] 7.21 1.31 0.661 0.678 0.247 0.289
STAGE [21] 6.52 1.11 0.732 0.747 0.365 0.139
SUICA (Ours) 5.66 0.85 0.797 0.792 0.447 0.343
Table 1: Quantitative benchmarking results on MOSTA dataset [2]. Bold figures are best scores and underlined figures are second-best.
Figure 4: Visual comparisons of predicted gene expressions on MOSTA dataset [2].
SUICA outperforms vanilla INR methods [48, 43] and STAGE [21] in predicting the gene expression of unseen spots in the MOSTA dataset [2] as presented in Tab. 1. SUICA not only achieves the best MAE, MSE and cosine similarity, exceeding the second-best method by 5.6%, but also achieves higher correlation, with a significant gap ahead by at least 4.5%. Apart from the reconstruction quality, SUICA surprisingly manages to strengthen the biological signals, by obtaining a bio-conservation score 5.4% higher than the runner-up method and even 3.9% above the ARI calculated with the ground truth.
Fig. 4 showcases SUICA’s ability to closely predict ground-truth gene expression levels, e.g., for RPLP0 gene that highly involves in the ribosomal function and consistently expresses across various regions of mouse embryo [50, 35], SUICA accurately predicts this uniform expression pattern, while other methods overemphasize it in specific regions. SUICA also accurately localizes AFP expression to the liver in the mouse embryo [19]. Additionally, SUICA is the only method to capture the expression of RP58, a transcription factor crucial for mouse brain development [34].
Beyond accurately predicting gene expression in unobserved regions, SUICA is also capable for imputation, enhancing the data to better reflect true underlying biological signatures. For example, despite its low expression level in the ground-truth, SUICA successfully imputes SEPT3, a gene involved in neuronal development, effectively restoring this signal in the brain region. A similar restoration is observed in ELAVL3 of the E10.5 embryo [8]. These results highlight SUICA’s capacity not only to interpolate but also to enrich underlying biological signatures, making it a useful tool for imputing and enhancing spatial gene expression data with high fidelity preserved.
Human and Mouse Brains Visium
Methods Visium-Human Brain Visium-Mouse Brain
@@
Nearest Interp. 5.36 0.824 0.781 0.365 5.45 0.870 0.736 0.566 0.292
FFN [48] 5.76 0.772 0.786 0.402 5.95 0.832 0.741 0.581 0.000587
SIREN [43] 6.58 0.756 0.747 0.196 5.35 0.878 0.804 0.647 0.359
STAGE [21] 6.19 0.795 0.772 0.223 4.55 0.918 0.825 0.666 0.140
SUICA (Ours) 4.99 0.860 0.846 0.445 3.68 0.932 0.800 0.660 0.393
Table 2: Quantitative benchmarking results on two Visium ST cases [36, 53]. Note that for Visium-Human Brain, there is no annotation of cell type for the evaluation of ARI. Bold figures are best scores and underlined figures are second-best.
Different from Stereo-seq datasets, Visium is a lower-resolution but more affordable ST technology. Tab. 2 shows that SUICA outperforms other methods on the Visium-Human Brain dataset [53]. Due to the lack of independent cell-type annotations, ARI is not applicable in this data. On the Visium-Mouse Brain dataset [36], SUICA demonstrates improved ARI, which indicates that the predictions can more accurately describe the underlying cellular heterogeneity.
Mouse hippocampus Slide-seqV2
Slide-seqV2 allows to sequence RNAs with a spatial resolution of 10 $mm$. The Mouse hippocampus dataset [44] is applied to evaluate the effectiveness of SUICA. As shown in Tab. 3, SUICA achieves a substantially lower MAE and a notably higher cosine similarity (marking a 16.5% improvement) compared to other methods. Fig. 5 illustrates the ground-truth and predicted gene expression from the benchmarking methods. Like other methods, SUICA accurately predicts the expression of hippocampus marker genes, such as HPCA (Hippocalcin) [39] and MALAT1, a gene abundantly expressed in the mouse brain [39]. SUICA also demonstrates its ability to impute the underlying biological signals, as evidenced by GABRB3, which is consistently and strongly expressed in the mouse hippocampus [47]. These findings suggest that SUICA provides reliable predictions that closely reflect ground-truth data, capturing key gene expression patterns in the Slide-seqV2 platforms.
Figure 5: Visual comparisons of the predicted marker gene expression patterns in the Slide-seqV2 Mouse hippocampus dataset [44].
[…]
Nearest Interp. 0.360 0.555 0.361 0.222 0.111
FFN [48] 0.378 0.499 0.442 0.274 0.0523
SIREN [43] 0.383 0.494 0.452 0.248 0.110
STAGE [21] 0.351 0.587 0.483 0.314 0.0361
SUICA (Ours) 0.265 0.752 0.473 0.308 0.111
Table 3: Quantitative results on Slide-seqV2 Mouse hippocampus dataset [44]. Bold figures are best scores and underlined figures are second-best.
@@ 4.4 Bio-conservation Analysis of Predicted Cell Type Clusters
Biological variance conservation (Bio-conservation) refers to the degree to which a computational method or model preserves biologically meaningful features, such as cell types, gene expression patterns, or cellular relationships, relative to the ground-truth data [26]. A high bio-conservation score indicates that the predicted data aligns closely with biological reality, capturing key aspects of cell identity, functional states, or tissue structure that are essential for accurate analysis.
In addition to quantitative bio-conservation metrics like ARI used in Section 4.3, Fig. 6 provides a visual assessment of bio-conservation by coloring predicted spots according to dominant cell types in individual clusters. For a fair comparison, we ensured that the number of clusters matched the number of cell-type labels in the dataset.
In the mouse embryo E16.5, apart from the highest ARI among benchmarking methods, SUICA also offers the closest spatial visualization of cell types to the ground truth. Compared to other methods, SUICA captures fine-grained structures, such as the choroid plexus in the brain, the gastrointestinal (GI) tract, and the muscle cells surrounding the heart. FFN’s [48] predictions appear chaotic, with most cell-type predictions misaligned with the ground-truth labels. While SIREN [43] correctly captures general cell types patterns, it overlooks many details detected by SUICA, such as the connective tissue around the cartilage primordium in the mouse’s foot region and the cavity surrounding the heart. STAGE [21] identifies several major cell types in the mouse embryo—such as liver and brain—but misses smaller, yet important, cell populations.
@@ 4.5 Ablation Study
We showcase two examples to illustrate how the proposed modules in SUICA contribute to the final model performance, namely the E16.5 embryo of Stereo-seq MOSTA (spatially dense, 121,756 cells, 26,159 genes, with FFN [48] as backbone) and the Human Brain Visium (spatially sparse, 4,910 cells, 36,592 genes, with SIREN [43] as backbone).
@@
Table 4: Ablation study on the key modules in SUICA.
As can be indicated from Tab. 4, the spatial density and sequencing depth may influence SUICA’s effectiveness. For the Embryo E16.5 dataset, the primary performance improvement comes from the AE and the quasi-classification loss. In contrast, in the Human Brain dataset, adding AE causes a performance decrease, while the GCN has a more significant impact on overall performance. Intuitively, we attribute this difference to the varying spatial sparsity of different ST techniques and GAE alleviates such issue by incorporating spatial context.
@@ 5 Conclusion
In this paper, we seek to model the super-high dimensional and sparse nature of ST data, enhancing both spatial resolution and gene expression with the smooth prior inherent in INRs. To this end, we present SUICA, a powerful INR variant tailored to model ST in a continuous and compact manner. Using a Graph Autoencoder, SUICA maps zero-inflated raw data into a lower-dimensional embedding space, preserving high-frequency details and making the complex embedding mapping more feasible for INRs. The INR fitted embeddings are then decoded to the raw expression with the decoding head preventing the prediction error from accumulating. To encourage the sparsity in the final predictions, we leverage a quasi-classification loss term as a regularizer, preserving both the sparsity and numerical fidelity of non-zero values. Extensive experiments on Stereo-seq, 10x Genomics Visium, and Slide-seqV2 datasets demonstrate SUICA’s effectiveness, yielding improvements in both in-silico metrics and biologically meaningful analyses, with enhanced spatial resolution and biological signatures. It is firmly believed by us that SUICA will be an inspiring attempt from both INRs and ST perspectives.
Supplementary Material
Due to the page limit of the main text, we have included some background information, some discussions regarding the model choices in SUICA, and some other experimental results we find meaningful for the readers.
6 Background

### v2

@@ 3.2.4 Decoding Head
A straightforward approach to decoding $\hat{\mathbf{z}}$ back to raw representations is to directly use the pre-trained GAE decoder for end-to-end INR tuning. However, this method encounters two primary challenges. (1) The embedding mapping of $\mathbf{x}\to\mathbf{z}$ is error-prone, and domain shifts can severely impair decoding performance, compounding with the inherent imperfections in GAE’s reconstruction. (2) A pre-trained decoder can become trapped in local minima, impeding the optimization of the embedding mapping if the INR depends on gradients from the decoder. To address (1)(2), we first warm up the INR with Equation 2 alone to stabilize the mapping, then attach the pre-trained decoder to learn the mapping from $\mathbf{z}\to\mathbf{y}$, leaving the INR fixed. This decoder-only training phase is designed to finetune a case-specific decoder that compensates for mapping errors and minimizes cumulative errors.
In INR regression tasks, norm-based loss functions (e.g., $\ell^{2}$ norm or MSE) are typically used, as they assume normally (Gaussian) distributed errors. However, for zero-inflated ST data, this assumption is invalid. To handle the imbalanced distribution of zero and non-zero values, we apply Dice Loss (Sudre et al., 2017), which is sensitive to class imbalance and treats the regression task as a quasi-classification one. Dice Loss optimizes for Intersection over Union (IoU) between the prediction map and binary ground truth. Specifically, we use the non-negative half of $\tanh(\cdot)$ to map network outputs into a pseudo-probability range $[0,1)$, and compute the intersection using element-wise Hadamard products with the ground truth. To avoid division by 0, Dice Loss is computed as:
@@ 4 Experiments
Figure 4: Visual comparisons of predicted marker gene expressions on MOSTA dataset (Chen et al., 2022a) and Slide-seqV2 Mouse hippocampus dataset (Stickels et al., 2021), with the descriptions of the markers attached below the results.
@@ 4.1 Datasets & Metrics
For a quantitative benchmarking, we involve a nanoscale resolution Stereo-seq (SpaTial Enhanced REsolution Omics-sequencing) dataset, MOSTA (Chen et al., 2022a). MOSTA consists of a total of 53 sagittal sections from C57BL/6 mouse embryos at 8 progressive stages using Stereo-seq, from which we take 1 slice for each stage (from E9.5 to E16.5) for benchmarking. In addition to Stereo-seq, we also leverage ST data by other common platforms, i.e., Slide-seqV2, 10x Genomics Visium (see Appendix) and MERFISH (see Appendix), to further demonstrate the generalization of SUICA.
As for evaluation of the fitting performance with super-high dimensional data, we focus on 3 aspects, namely numerical fidelity, statistical correlation, and bio-conservation. For numerical fidelity, we apply MSE, MAE (Mean Absolute Error) and cosine similarity to measure the significant, subtle and directional errors between the predicted and ground-truth values respectively. Note that we only measure numerical fidelity on non-zero values considering the zero-inflated distribution of ST. To measure the statistical correlation, we employ Pearson Correlation Coefficient and Spearman’s Rank Correlation Coefficient (Spearman’s $\rho$), with both of them ranging from -1 to 1. Lastly, to evaluate how well the prediction preserves cellular heterogeneity and spatial coherence within the microenvironment of the slice, we use the Adjusted Rand Index (ARI) as a metric for quantifying bio-conservation leveraging the independent hand-crafted cell type annotations.
4.2 Evaluation Protocol
As is the conventional data flow for INRs, SUICA infers gene expressions with coordinates as inputs after fitting the given $(\mathbf{x},\mathbf{y})$ pairs. We apply SUICA to perform degradation-agnostic reconstruction upon ST data under various common degradations, including spatial sparsity, gene drop-out, and noise, to which we refer as spatial imputation, gene imputation and denoising respectively for clarity. Accordingly, we follow different evaluation protocols: for spatial imputation, we randomly sample 80% of the spots for training, leaving the rest 20% for evaluation; for gene imputation, we randomly mute 70% of the elements in the data matrices; for denoising, a standard Gaussian noise is injected to the raw data. Note that, such reconstruction is completely based on the great approximation power and internal smoothness of INRs without any beforehand knowledge of the degradation type in a reference-free manner.
For comparison, we compare SUICA with rule-based well-known INR variants (FFN (Tancik et al., 2020) and SIREN (Sitzmann et al., 2020)), and the SOTA learning-based reference-free baseline model STAGE (Li et al., 2024).
@@ 4.3 Implemented Details
We implement SUICA with PyTorch. To construct the $k$-NN graph for GCN, we set $k=5$, including the given cell itself. During the pre-training phase of GAE, we use Adam with a learning rate of $1e^{-5}$ for 200 epochs. After obtaining low-dimensional cell embeddings, we train the INR with Adam at a learning rate of $1e^{-4}$ for $1k$ epochs to learn the embedding mapping. Subsequently, the INR is frozen, with the pre-trained decoder trained for an additional $1k$ epochs using Adam with the same learning rate. STAGE (Li et al., 2024) is trained till convergence for benchmarking. All experiments can be conducted on 1 NVIDIA RTX 4090.
4.4 Spatial Imputation
@@
Table 1: Quantitative benchmarking results of spatial imputation on MOSTA dataset (Chen et al., 2022a) and Slide-seqV2 Mouse hippocampus dataset (Stickels et al., 2021). Bold figures are best scores and underlined figures are second-best. The respective reference ARI scores are 0.312 (Stereo-seq MOSTA) and 0.182 (Mouse hippocampus Slide-seqV2). MAE/MSE: ×$\times 10^{-2}$ for Stereo-seq MOSTA.
Methods Stereo-seq MOSTA Mouse hippocampus Slide-seqV2
@@
FFN (Tancik et al., 2020) 6.51 1.20 0.706 0.718 0.400 0.143 0.378 0.215 0.499 0.442 0.274 0.0523
SIREN (Sitzmann et al., 2020) 7.21 1.31 0.661 0.678 0.247 0.289 0.383 0.216 0.494 0.452 0.248 0.110
STAGE (Li et al., 2024) 6.52 1.11 0.732 0.747 0.365 0.139 0.351 0.198 0.587 0.483 0.314 0.0361
SUICA (Ours) 5.66 0.85 0.797 0.792 0.447 0.343 0.265 0.125 0.752 0.473 0.308 0.111
SUICA outperforms INR methods (Tancik et al., 2020; Sitzmann et al., 2020) and STAGE (Li et al., 2024) in predicting the gene expression of unseen spots in the MOSTA dataset (Chen et al., 2022a) as presented in Table 1. SUICA not only achieves the best MAE, MSE and cosine similarity, but also achieves higher correlation. Besides, SUICA surprisingly manages to strengthen the biological signals, even 3.9% above the ARI calculated with the ground truth.
Figure 4 showcases SUICA’s ability to closely predict ground-truth gene expression levels, e.g., for RPLP0 gene that highly involves in the ribosomal function and consistently expresses across various regions of mouse embryo (Taylor & Pikó, 1992; Ozadam et al., 2023), SUICA accurately predicts this uniform expression pattern, while other methods overemphasize it in specific regions. SUICA also accurately localizes AFP expression to the liver in the mouse embryo (Kwon et al., 2006). Beyond accurately predicting gene expression in unobserved regions, SUICA is also capable for imputation, enhancing the data to better reflect true underlying biological signatures. For example, despite its low expression level in the ground-truth, SUICA successfully imputes SEPT3, a gene involved in neuronal development, effectively restoring this signal in the brain region. These results highlight SUICA’s capacity not only to interpolate but also to enrich underlying biological signatures, making it a useful tool for imputing and enhancing spatial gene expression data with high fidelity preserved.
Mouse hippocampus Slide-seqV2
Slide-seqV2 allows to sequence RNAs with a spatial resolution of 10 $mm$. The Mouse hippocampus dataset (Stickels et al., 2021) is applied to evaluate the effectiveness of SUICA. As shown in Table 1, SUICA achieves a substantially lower MAE and a notably higher cosine similarity (marking a 16.5% improvement) compared to other methods. Figure 4 illustrates the ground-truth and predicted gene expression from the benchmarking methods. Like other methods, SUICA accurately predicts the expression of hippocampus marker genes, such as HPCA (Hippocalcin) (Park et al., 2017) and MALAT1, a gene abundantly expressed in the mouse brain (Park et al., 2017). SUICA also demonstrates its ability to impute the underlying biological signals, as evidenced by GABRB3, which is consistently and strongly expressed in the mouse hippocampus (Tanaka et al., 2012). These findings suggest that SUICA provides reliable predictions that closely reflect ground-truth data, capturing key gene expression patterns in the Slide-seqV2 platforms.
Table 2: Quantitative benchmarking results of gene imputation and denoising on MOSTA dataset (Chen et al., 2022a). Bold figures are best scores and underlined figures are second-best. MAE/MSE: ×$\times 10^{-2}$.
Methods Gene Imputation Denoising
@@
FFN (Tancik et al., 2020) 4.88 0.963 0.731 0.610 0.251 7.90 1.95 0.266 0.285 0.0523
SIREN (Sitzmann et al., 2020) 6.44 1.12 0.675 0.652 0.124 7.91 1.97 0.112 0.103 0.0166
STAGE (Li et al., 2024) 4.69 0.738 0.802 0.705 0.264 7.60 1.66 0.606 0.630 0.182
SUICA (Ours) 4.30 0.724 0.798 0.714 0.317 6.03 0.934 0.733 0.737 0.379
@@ 4.5 Bio-conservation Analysis of Predicted Cell Type Clusters
Biological variance conservation (Bio-conservation) refers to the degree to which a computational method or model preserves biologically meaningful features, such as cell types, gene expression patterns, or cellular relationships, relative to the ground-truth data (Luecken et al., 2022). A high bio-conservation score indicates that the predicted data aligns closely with biological reality, capturing key aspects of cell identity, functional states, or tissue structure that are essential for accurate analysis.
In addition to quantitative bio-conservation metrics like ARI, Figure 5 provides a visual assessment of bio-conservation by coloring predicted spots according to dominant cell types in individual clusters. For a fair comparison, we ensured that the number of clusters matched the number of cell-type labels in the dataset.
In E16.5, apart from the highest ARI among benchmarking methods, SUICA also offers the closest spatial visualization of cell types to the ground truth. Compared to other methods, SUICA captures fine-grained structures, such as the choroid plexus in the brain, the gastrointestinal (GI) tract, and the muscle cells surrounding the heart. FFN’s (Tancik et al., 2020) predictions appear chaotic, with most cell-type predictions misaligned with the ground-truth labels. While SIREN (Sitzmann et al., 2020) correctly captures general cell types patterns, it overlooks many details detected by SUICA, such as the connective tissue around the cartilage primordium in the mouse’s foot region and the cavity surrounding the heart. STAGE (Li et al., 2024) identifies several major cell types in the mouse embryo—such as liver and brain—but misses smaller, yet important, cell populations.
@@
4.6 Gene Imputation & Denoising
Since INRs are known for its internal smoothness, SUICA is able to perform channel-wise gene imputation and denoising upon muted or contaminated gene expressions. In this way, SUICA can be seen as a reference-free degradation-agnostic restoration method for ST.
We benchmark the quantitative results in Table 2 upon Stereo-seq MOSTA (Chen et al., 2022a). Note that the pipeline is exactly identical except for the degradation pattern. SUICA manages to handle the contaminated inputs in almost all metrics. It also shows that GAE manages to make it easier for INRs to model the smoothness in the super-high dimensional space of gene expressions.
@@ 4.7 Ablation Study
We showcase two examples to illustrate how the proposed modules in SUICA contribute to the final model performance, namely the E16.5 embryo of Stereo-seq MOSTA (spatially dense, 121,756 cells, 26,159 genes, with FFN (Tancik et al., 2020) as backbone) and the Visium Human Brain (spatially sparse, 4,910 cells, 36,592 genes, with SIREN (Sitzmann et al., 2020) as backbone), under the setting of spatial imputation.
Table 3: Ablation study on model design of SUICA. MSE: ×$\times 10^{-2}$ for Embryo E16.5 while ×$\times 10^{-3}$ for Human Brain.
@@
As can be indicated from Table 3, the spatial density and sequencing depth may influence SUICA’s effectiveness. For the Embryo E16.5 dataset, the primary performance improvement comes from the AE and the quasi-classification loss. In contrast, in the Human Brain dataset, adding AE causes a performance decrease, while the GCN has a more significant impact on overall performance. Intuitively, we attribute this difference to the varying spatial sparsity of different ST techniques and GAE alleviates such issue by incorporating spatial context.
@@ 5 Conclusion
In this paper, we seek to model the super-high dimensional and sparse nature of ST data, enhancing both spatial resolution and gene expression with the smooth prior inherent in INRs. To this end, we present SUICA, a powerful INR variant tailored to model ST in a continuous and compact manner. Using a Graph Autoencoder, SUICA maps zero-inflated raw data into a lower-dimensional embedding space, preserving high-frequency details and making the complex embedding mapping more feasible for INRs. The INR fitted embeddings are then decoded to the raw expression with the decoding head preventing the prediction error from accumulating. To encourage the sparsity in the final predictions, we leverage a quasi-classification loss term as a regularizer, preserving both the sparsity and numerical fidelity of non-zero values. Extensive experiments on Stereo-seq, 10x Genomics Visium, and Slide-seqV2 datasets demonstrate SUICA’s effectiveness, yielding improvements in both in-silico metrics and biologically meaningful analyses, with enhanced spatial resolution and biological signatures. It is firmly believed by us that SUICA will be an inspiring attempt from both INRs and ST perspectives. We encourage readers to check the appendix for complementary insights and extended evaluations.
Acknowledgements
This work was supported in part by Three Lakes Foundation, the Canadian Institutes of Health Research (CIHR) [PJT-180505], the Fonds de recherche du Québec - Santé (FRQS) [295298, 295299], the Meakins-Christie Chair in Respiratory Research, AMED [JP23tm0524003], JSPS KAKENHI [24KK0209, 24K22318, 22H00529], and JST-Mirai Program [JPMJMI23G1].
Impact Statement
SUICA offers a comprehensive solution to the challenges of spatial sparsity, gene drop-out, and noise in spatial transcriptomics data analysis. It is compatible with a wide range of sequencing platforms, making it broadly accessible to research teams. Given the high cost of obtaining ST data, SUICA empowers the scientific community to enhance data quality and accelerate the discovery of novel biological insights.
Appendix A Background